Plant I of a manufacturing organisation employs 5 production and 3 maintenance engineers; another plant II of same organisation employs 4 production and 5 maintenance engineers. From any one of these plants, a single selection of two engineers is made. Find the probability that one them would be production engineer and the other maintenance engineer.
1
Expert's answer
2020-01-20T09:30:29-0500

As we should choose from any of two plants, the probability that the choice will be made at the first plant is "P1=\frac{1}{2}" , and at the second plant is "P2=\frac{1}{2}". If we have plant I the probability to collect different two engineers is a summation of two successive events. The first event is a product of two successive events: we initially choose maintenance engineer "P1_m=\frac{3}{8}" (at the first plant worked 8 engineers and only 3 maintenance) and then the production engineer"P1_p=\frac{5}{7}" (at the first plant after first choice remained 7 engineers and 5 production engineers). The second event is 'we initially get production engineer' with the probability "P1^{'}_p= \frac{5}{8}" and then some maintenance engineer with the probability "P1^{'}_m=\frac{ 3}{7}". The overall probability for a right selection in the first plant will be "P1_{m+p}=P1_m \cdot P1_p+P1^{'}_p\cdot P1^{'}_m=\frac {3}{8}\cdot \frac {5}{7}+\frac {5}{8}\cdot \frac {3}{7}=\frac {15}{28}" .

Similarly for second plant we get

"P2_{m+p}=P2_m\cdot P2_p+P2^{'}_p\cdot P2^{'}_m=\frac {5}{9}\cdot \frac {4}{8}+\frac {4}{9}\cdot \frac {5}{8}=\frac {5}{9}"

The probability of choosing different engineers in the organization will be

"P_{m+p}=P1\cdot P1_{m+p}+P2\cdot P2_{m+p}=\frac {1}{2}\cdot \frac {15}{28}+\frac {1}{2}\cdot \frac {5}{9}=\frac {1}{2}\cdot \frac {135+140}{252}=\frac{275}{504}"

Answer: The probability that one of two engineers will be production engineer and the other maintenance engineer is "\frac{275}{504}" .